OpenStudy (anonymous):

What mixed number falls in between 5 and 31/6

sammixboo (sammixboo):

Okay now we need to change the fraction \(\sf \large \color{magenta}{\frac{31}{6}}\) into a mixed fraction... \(\sf \large \color{magenta}{\frac{31}{6}}\) ---- \(\sf \large \color{magenta}{5\frac{1}{6}}\)

sammixboo (sammixboo):

Since there is not fraction between \(\sf \large \color{magenta}{5}\) and \(\sf \large \color{magenta}{5\frac{1}{6}}\) we need to take the fraction \(\sf \large \color{magenta}{5\frac{1}{6}}\) and multiply it by two... \(\sf \large \color{magenta}{5\frac{1 x 2}{6 x 2}}\) ---- \(\sf \large \color{magenta}{5\frac{2}{12}}\)

sammixboo (sammixboo):

What fraction is in between \(\sf \large \color{magenta}{5}\) and \(\sf \large \color{magenta}{5\frac{2}{12}}\)